OpenStudy (opentowherelifeis):

Evaluate the expression when x=-1 and y = 4 9x ^-5 y^-2 ___________ y

OpenStudy (opentowherelifeis):

@satellite73

OpenStudy (anonymous):

\[\frac{ 9x^{-5}y^{-2} }{ y }\] like this?

OpenStudy (opentowherelifeis):

yes

OpenStudy (anonymous):

example..\[y^{-3}=\frac{ 1 }{ y^3 }\]

OpenStudy (anonymous):

\[\frac{ 9x^{-5}y^{-2} }{ y }=\frac{ 9 }{ x^5y^{2+1} }\]

OpenStudy (anonymous):

\[=\frac{ 9 }{ x^5y^3 }\]Then,substiute x=-1 and y=4 into the expression

OpenStudy (anonymous):

\[=\frac{ 9 }{ (-1)^5\times(4)^3 }\]

OpenStudy (anonymous):

so what is the value for\[(-1)^5\]?@Opentowherelifeis

OpenStudy (opentowherelifeis):

would it be - 9 over 64

OpenStudy (anonymous):

correct!

OpenStudy (opentowherelifeis):

yay . THANKS FOR THE HELP :) here is your medal

OpenStudy (anonymous):

(-1)^5 gives u -1 and (4)^3 gives u 64 \[=\frac{ 9 }{ -1\times64 }\]\[=-\frac{ 9 }{ 64 }\]

OpenStudy (anonymous):

np :)